prove that xcosecx=1+(x)^2/6+(7/360)*x^4


1
Expert's answer
2022-01-28T03:44:53-0500

The Maclaurin series for "\sin x" is


"\sin x=x-\dfrac{x^3}{3!}+\dfrac{x^5}{5!}-\dfrac{x^7}{7!}+..."

Then


"x\cosec x=\dfrac{x}{\sin x}=\dfrac{x}{x-\dfrac{x^3}{3!}+\dfrac{x^5}{5!}-\dfrac{x^7}{7!}+...}"

"=\dfrac{1}{1-(\dfrac{x^2}{3!}-\dfrac{x^4}{5!}+\dfrac{x^6}{7!}+...)}"

The Maclaurin series for "\dfrac{1}{1-y}" is


"\dfrac{1}{1-y}=1+y+y^2+y^3+..., -1<y<1"

Then


"\dfrac{1}{1-(\dfrac{x^2}{3!}-\dfrac{x^4}{5!}+\dfrac{x^6}{7!}+...)}""=1+(\dfrac{x^2}{3!}-\dfrac{x^4}{5!}+\dfrac{x^6}{7!}+...)"

"+(\dfrac{x^2}{3!}-\dfrac{x^4}{5!}+\dfrac{x^6}{7!}+...)^2"

"+(\dfrac{x^2}{3!}-\dfrac{x^4}{5!}+\dfrac{x^6}{7!}+...)^3+..."

"=1+\dfrac{x^2}{3!}-\dfrac{x^4}{5!}+\dfrac{x^4}{3!3!}+\dfrac{x^6}{7!}-\dfrac{2x^6}{3!5!}+\dfrac{x^6}{3!3!3!}+..."

"=1+\dfrac{x^2}{6}+\dfrac{7x^4}{360}+..."

Therefore


"x\cosec x=1+\dfrac{x^2}{6}+\dfrac{7x^4}{360}+..."